# Visualizing LLM Latent Space Geometry Through Dimensionality Reduction

### v1

@@ Abstract
Large language models (LLMs) achieve state-of-the-art results across many natural language tasks, but their internal mechanisms remain difficult to interpret. In this work, we extract, process, and visualize latent state geometries in Transformer-based language models through dimensionality reduction. We capture layerwise activations at multiple points within Transformer blocks and enable systematic analysis through Principal Component Analysis (PCA) and Uniform Manifold Approximation (UMAP). We demonstrate experiments on GPT-2 and LLaMa models, where we uncover interesting geometric patterns in latent space. Notably, we identify a clear separation between attention and MLP component outputs across intermediate layers, a pattern not documented in prior work to our knowledge. We also characterize the high norm of latent states at the initial sequence position and visualize the layerwise evolution of latent states. Additionally, we demonstrate the high-dimensional helical structure of GPT-2’s positional embeddings, the sequence-wise geometric patterns in LLaMa, and experiment with repeating token sequences. We aim to support systematic analysis of Transformer internals with the goal of enabling further reproducible interpretability research. We make our code available at https://github.com/Vainateya/Feature_Geometry_Visualization.
@@ 2.1 Transformers
Originally introduced by Vaswani et al. in 2017 (vaswani2023attentionneed), Transformers have achieved great notoriety for their state-of-the-art performance across most language modeling tasks. Despite this success, their internal mechanisms remain difficult to understand, motivating extensive work in mechanistic interpretability (see Section 1). Since our experiments focus on analyzing decoder-only Transformers (e.g., GPT-2 (radford2019language) and LLaMa (touvron2023llamaopenefficientfoundation)), we begin by providing a brief overview of the Transformer architecture. To ensure clarity amongst various terminologies in recent literature, we also establish a consistent set of terms that will be used throughout this paper. That is, we formalize a decoder-only Transformer as a sequence of blocks, each consisting of four primary components/layers: the normalization layer (e.g., layer norm or RMSNorm) preceding the attention component, the multi-head self-attention, the normalization layer preceding the MLP, and the MLP (multilayer perceptron). Although the original Transformer architecture utilized a post-norm design, where the normalization layer came after the attention or MLP component, most modern architectures (GPT-2 and LLaMa) adopt a pre-norm design. Thus, we depict the pre-norm architecture in figures 1 and 3. We use the term “final norm” to refer to the normalization applied after the last Transformer block. This is immediately followed by the unembedding layer, which maps the final latent representation into vocabulary logits for token prediction. Finally, we use 0-based indexing when referring to blocks, layers, and sequence positions throughout the paper. For example, “block 0” denotes the first block in the Transformer, and “sequence position 0” denotes the first token position in a sequence. In order to avoid ambiguity between “first” and index “1”, we will refer to the “first” (0-th) item as the “initial” item. While Transformers are often conceptualized as a sequential operation through each of their components with residual connections between layers, we adopt the mathematically equivalent perspective that these skip connections collectively form the central communication channel through which all components interact. We refer to this pathway as the residual stream. (elhage2021mathematical).
@@
Figure 1 illustrates these differing perspectives. In this residual stream-centric view, all components of a Transformer (the token embedding, attention heads, MLP layers, and unembedding layer) communicate with each other by reading and writing to different subspaces of the residual stream. Rather than thinking of information flowing sequentially through layers, we conceptualize each layer as reading its input from the residual stream (by performing a linear projection), and then writing its result to the residual stream by adding a linear projection back in (elhage2021mathematical). This additive structure preserves a path for an identity pathway to flow through the model and into each component, unobstructed by any operation other than the direct sum of component updates back into the residual stream. We highlight the importance of the residual stream perspective for two reasons. Firstly, it allows us to treat all components of the Transformer as operating within a shared representational space, intuitively describing the collaborative nature of component interactions. This shared space enables different components to develop a collective understanding of features and their corresponding directions, making it possible to interpret representational changes as coordinated rather than isolated transformations. Secondly, this framing provides a conceptual basis for analyzing how the geometry of these shared representations evolves throughout the residual stream, thereby offering insights into how components jointly shape the model’s feature space.
@@ 2.2 Linear Representation Hypothesis
LLMs learn internal features, which are abstract properties or characteristics of data that guide their predictions and generalization. In language modeling, such features often correspond to linguistic attributes such as gender, tense, or sentiment. Mikolov et al. (mikolov-etal-2013-linguistic) showed that word embeddings capture these features through simple vector arithmetic. For example, $\langle\text{King}\rangle-\langle\text{Man}\rangle+\langle\text{Woman}\rangle\approx\langle\text{Queen}\rangle$, where angle brackets denote embedding vectors. This provided the first evidence that semantic and syntactic relations can be represented as linear patterns in latent space. This empirical finding has since been formalized in the context of large language models as the Linear Representation Hypothesis (LRH) (park2024linearrepresentationhypothesisgeometry), which posits that high-level concepts are represented as directions (i.e., one-dimensional subspaces) in representation space. This implies an intuition where learned representations are approximately linear combinations of feature directions.
[…]
@@ 2.3 Dimensionality Reduction
Modern neural networks generate latent representations that live in extremely high-dimensional spaces. While these representations contain rich structural information, they are difficult to interpret directly. Dimensionality reduction techniques serve as a bridge between high-dimensional representations and human visualization. The goal is to compress embeddings into lower dimensions while preserving the salient structure, e.g., variance, clusters, or manifold geometry, so that latent state behavior can be more easily examined. In this work, we study both Principal Component Analysis (PCA) and Uniform Manifold Approximation (UMAP) as the primary methods for dimensionality reduction and visualization.
PCA is a widely used method for linear dimensionality reduction. PCA operates by finding a new orthogonal basis that captures directions of maximal variance in the data (shlens2014tutorialprincipalcomponentanalysis). By projecting embeddings onto these principal axes, PCA allows one to rank-order components by their explanatory power. The method is computationally efficient, analytically grounded in linear algebra through eigenvalue decomposition or singular value decomposition, and does not require hyperparameter tuning. However, PCA is limited to capturing linear relationships and assumes that variance corresponds to meaningful structure, an assumption that may fail in non-Gaussian or highly nonlinear datasets (shlens2014tutorialprincipalcomponentanalysis). Despite these caveats, PCA remains a powerful baseline, is widely used in our experiments, and provides a stable reference frame for comparing geometric patterns.
On the other hand, Uniform Manifold Approximation and Projection (UMAP) (mcinnes2020umapuniformmanifoldapproximation), provides a nonlinear alternative designed to preserve both local neighborhoods and some aspects of global manifold structure. Rather than identifying variance-maximizing axes, UMAP builds a graph-based representation of data as a topological manifold and then optimizes a low-dimensional embedding that preserves this structure. In practice, UMAP is effective at revealing clustering patterns in latent states, producing interpretable visualizations. We chose UMAP over other nonlinear alternatives such as t-SNE (JMLR:v9:vandermaaten08a) because it better maintains global relationships while still revealing local clustering patterns. Its flexibility comes at the cost of tunable hyperparameters (e.g., number of neighbors, minimum distance) and a lack of linear structure, but it has become a widely adopted tool in machine learning for exploring latent structure.
@@ 2.4 Positional Embeddings/Encodings
In many early Transformer architectures, such as GPT-2 (used in our experiments), positional information is injected via learned positional embeddings. Each token position in the input sequence is associated with a learned vector that is added to its corresponding token embedding before entering the first Transformer block. Because these embeddings are learned during training, they often exhibit consistent geometric patterns that reflect the model’s internal representation of positional order. Within the context of our work, learned embeddings provide a straightforward entry point for feature geometry analysis, serving as a baseline for how position-dependent features evolve through the residual stream.
On the other hand, the LLaMa model we use employs a specific form of functional positional encoding called Rotary Positional Encodings (RoPE). Rather than adding fixed position vectors, RoPE applies a position-dependent rotation to query and key vectors within each attention head. This rotation encodes relative position information directly into the attention mechanism, allowing the model to generalize more effectively to longer sequences. Conceptually, RoPE embeds position in the geometry of relationships between tokens, rather than as an explicit positional vector.
This distinction serves as the key motivation behind Effects of Sequence Position experiments in section 4.5 and Repeating Token Experiments in section 4.6.
@@ 3.1 Pipeline
The pipeline supports two modes of operation. In the default “text” mode, text passages are sampled from a text dataset. For our experiments, we ran the model through the well-known Project Gutenberg (PG-19) dataset (rae2019compressivetransformerslongrangesequence), consisting of a corpus of books written before 1919. Each passage is tokenized and truncated or padded to a fixed sequence length, ensuring uniformity across samples. In “singular” mode, by contrast, the pipeline probes individual tokens directly by iterating over the model’s vocabulary. This allows for fine-grained analysis of token-level representations without contextual interference. Both modes are parameterized by the number of samples and the desired sequence length (set at $1$ for the singular mode).
Figure 3 shows each of the capture points along the Transformer architecture. For each block, we record the outputs of each layer: both normalizations, the attention module, and the MLP. We denote the raw outputs of the attention and MLP components and their preceding norms as “pre-add” (points 1, 2, 4, and 5 in Figure 3). From these activations, we also compute the residual stream after each addition (post-attention and post-MLP), capturing the points where new computations are integrated into the residual pathway. We denote these as “post-add” (points 3, 6 in Figure 3). In total, this yields six distinct captures per Transformer block.
@@ 3.2 Dimensionality Reduction and Visualization
Captured latent states are visualized in 22D following dimensionality reduction. Not all captured latent states must be used for dimensionality reduction and visualization. Instead, a subset may be selected (e.g., only latent states from particular layers). For dimensionality reduction, both PCA and UMAP are supported, with a GPU-accelerated implementation for UMAP (via cuML (raschka2020machinelearningpythonmain)) utilized for visualizing large amounts of latent states. Before dimensionality reduction is performed, augmentations to the latent states can be applied. These include:
@@
  • Averaging the latent state vectors over any dimension(s) (sample dimension, sequence dimension, layer dimension). This may be done to reduce the variance of a certain dimension to concentrate on the structure present in another.
PCA transform is performed without mean centering to preserve the position of the origin.
We utilize a customizable color-coding scheme to distinguish the layer and sequence dimensions. In our experiments, attention layers are blue while MLP layers are red. Darker shades indicate earlier layers while lighter shades indicate later ones. For the sequence dimension, we use a gradient to indicate earlier vs. later positions.
Fitted dimensionality reduction models can be reused; PCA fitted on one set of latent states can be applied to a different set of latent states, and visualizations can preserve x- and y-axis limits between runs. This is especially useful for ablation studies, where subtle differences between experimental conditions must be compared in a consistent coordinate frame. Additionally, dimensionality reduction can be performed, which reduces the number of dimensions to a number greater than $2$. Then, any pair of the reduced dimensions can be visualized together in 22D (e.g., dimensions $2$ and $5$ or $3$ and $7$). Figure 11 shows the result of visualizing all possible pairs of dimensions after dimensionality reduction to $6$ dimensions (totaling ${6\choose 2}=15$ pairs), then stitching the resulting visualizations into a grid.
@@ 4.1 Experimental Setup
We used GPT-2 Large (774774M) and LLaMa-7B to generate the latent states. On all results using the PG-19 dataset, we use the following input shapes:
@@
We note that we use the maximum input length possible for both models. LLaMa is trained with a <BOS> (beginning of sentence) token prepended to all inputs. As a result, we prepend the <BOS> token to the beginning of all PG-19 inputs to LLaMa. After collecting the LLaMa latent states, we reduced the dimensionality of the latent states from $4096$ to $512$ using PCA. These dimensionality-reduced latent states were subsequently used in all visualizations except those in section 4.6. They were not used in any analyses of latent state norms; those used the original full dimensionality latent states for accurate results. All GPT-2 latent states were used in their full dimensionality.
Existing research demonstrates how different blocks in a Transformer have different roles. While earlier blocks focus on converting tokens into concepts (nanda2023factfinding) and the final blocks heavily denoise and refine the output, intermediate (middle) blocks, which are more robust to swapping and deletion, gradually create and refine intermediate features in a shared representation space (sun2025transformerlayerspainters; lad2025remarkablerobustnessllmsstages). For some of our experiments, we are interested in investigating the behavior of the intermediate block latent states without interference from earlier or later blocks. As a result, we define here the specific “intermediate” blocks/layers which we will later refer to in our results and discussions for both our tested models. Because there does not exist a clear methodology to define concrete boundaries for the "intermediate" blocks, the definitions we make are somewhat arbitrary. Nevertheless, it is important for us to have consistent definitions across our analyses. For GPT-2, we define these as the layers in blocks 22-88 (from a range of 0-$11$), while for LLaMa, they are the layers in blocks 66-$27$ (from a range of 0-$31$).

### v2

@@ Abstract
Large language models (LLMs) achieve state-of-the-art results across many natural language tasks, but their internal mechanisms remain difficult to interpret. In this work, we extract, process, and visualize latent state geometries in Transformer-based language models through dimensionality reduction. We capture layerwise activations at multiple points within Transformer blocks and enable systematic analysis through Principal Component Analysis (PCA) and Uniform Manifold Approximation and Projection (UMAP). We demonstrate experiments on GPT-2 and LLaMa models, where we uncover interesting geometric patterns in latent space. Notably, we identify a clear separation between attention and MLP component outputs across intermediate layers, a pattern not documented in prior work to our knowledge. We also characterize the high norm of latent states at the initial sequence position and visualize the layerwise evolution of latent states. Additionally, we demonstrate the high-dimensional helical structure of GPT-2’s positional embeddings and the sequence-wise geometric patterns in LLaMa. We make our code available at https://github.com/Vainateya/Feature_Geometry_Visualization.
@@ 2.1 Transformers
Originally introduced by Vaswani et al. in 2017 (vaswani2023attentionneed), Transformers have achieved great recognition for their state-of-the-art performance across most language modeling tasks. Despite this success, their internal mechanisms remain difficult to understand, motivating extensive work in mechanistic interpretability (see Section 1). Since our experiments focus on analyzing decoder-only Transformers (e.g., GPT-2 (radford2019language) and LLaMa (touvron2023llamaopenefficientfoundation)), we begin by providing a brief overview of the Transformer architecture. To ensure clarity amongst various terminologies in recent literature, we also establish a consistent set of terms that will be used throughout this paper. That is, we formalize a decoder-only Transformer as a sequence of blocks, each consisting of four primary components/layers: the normalization layer (e.g., layer norm or RMSNorm) preceding the attention component, the multi-head self-attention, the normalization layer preceding the MLP (multilayer perceptron), and the MLP. Although the original Transformer architecture utilized a post-norm design, where the normalization layer comes after the attention or MLP component, most modern architectures (e.g. GPT-2 and LLaMa) adopt a pre-norm design. Thus, we depict the pre-norm architecture in figures 1 and 3. We use the term “final norm” to refer to the normalization applied after the last Transformer block. This is immediately followed by the unembedding layer, which maps the final latent representation into vocabulary logits for token prediction. Finally, we use 0-based indexing when referring to blocks, layers, and sequence positions throughout the paper. For example, “block 0” denotes the first block in the Transformer, and “sequence position 0” denotes the first token position in a sequence. In order to avoid ambiguity between “first” and index “1”, we will refer to the “first” (0-th) item as the “initial” item. While Transformers are often conceptualized as a sequential operation through each of their components with residual connections between layers, we adopt the mathematically equivalent perspective that these skip connections collectively form the central communication channel through which all components interact. We refer to this pathway as the residual stream, as formalized by Elhage et al. in 2021 (elhage2021mathematical).
@@
Figure 1 illustrates these differing perspectives. In this residual stream-centric view, all components of a Transformer (the token embedding, attention heads, MLP layers, and unembedding layer) communicate with each other by reading and writing to different subspaces of the residual stream. Rather than thinking of information flowing sequentially through layers, we conceptualize each layer as reading its input from the residual stream (by performing a linear projection), and then writing its result to the residual stream by adding a linear projection back in (elhage2021mathematical). This additive structure preserves a path for an identity pathway to flow through the model and into each component, unobstructed by any operation other than the direct sum of component updates back into the residual stream. We highlight the importance of the residual stream perspective for two reasons. Firstly, it allows us to treat all components of the Transformer as operating within a shared representation space, intuitively illustrating the collaborative nature of component interactions. This shared space enables different components to develop a collective understanding of features and their corresponding directions, making it possible to interpret representational changes as coordinated rather than isolated transformations. Secondly, this framing provides a conceptual basis for analyzing how the geometry of these shared representations evolves throughout the residual stream, thereby offering insights into how components jointly shape the model’s feature space.
@@ 2.2 Linear Representation Hypothesis
LLMs learn internal features, which are abstract properties or characteristics of data that guide their predictions and generalization. In language modeling, such features often correspond to linguistic attributes such as gender, tense, or sentiment. Mikolov et al. in 2013 (mikolov-etal-2013-linguistic) showed that word embeddings capture these features through simple vector arithmetic. For example, $\langle\text{King}\rangle-\langle\text{Man}\rangle+\langle\text{Woman}\rangle\approx\langle\text{Queen}\rangle$, where angle brackets denote embedding vectors. This provided the first evidence that semantic and syntactic relations can be represented as linear patterns in latent space. This empirical finding has since been formalized in the context of large language models as the Linear Representation Hypothesis (LRH) (park2024linearrepresentationhypothesisgeometry), which posits that high-level concepts are represented as directions (i.e., one-dimensional subspaces) in representation space. This implies an intuition where learned representations are approximately linear combinations of feature directions.
LRH is significant for mechanistic interpretability in two respects. First, LRH implies a property of decomposability, which conveniently states that complex representations can be analyzed in terms of independent, concept-aligned features. This property provides a tractable foundation for understanding LLM latent space dynamics. Second, under the assumption that features are represented linearly, dimensionality reduction methods can more reliably reveal the organization and combination of features within the latent space. In this framing, each feature $i$ may be represented by a unit vector $e_{i}$, with an input’s representation expressed as $\sum x_{i}e_{i}$, where $x_{i}$ denotes the feature’s strength or presence (elhage2022superposition). With that said, Linear Representation Hypothesis remains a hypothesis, and there are limitations to the interpretations it can provide. Transformer models implement many behaviors which are extremely complex, not yet understood, and highly non-linear. Engels et al. in 2025 engels2025languagemodelfeaturesonedimensionally explored features which are not one-dimensionally linear. Additionally, the Superposition Hypothesis (elhage2022superposition) posits that in order to represent a large number of independent features in a space with limited dimensionality, neural networks permit a certain amount of interference between features by embedding them into almost-orthogonal directions instead of fully orthogonal directions. Although still linear, this degree of interference potentially clouds the benefits of interpretability and decomposability offered by LRH.
@@ 2.3 Dimensionality Reduction
Modern neural networks generate latent representations that live in extremely high-dimensional spaces. While these representations contain rich structural information, they are difficult to interpret directly. Dimensionality reduction techniques serve as a bridge between high-dimensional representations and human visualization. The goal is to compress embeddings into lower dimensions while preserving the salient structure, e.g., variance, clusters, or manifold geometry, so that latent state behavior can be more easily examined. In this work, we study both Principal Component Analysis (PCA) and Uniform Manifold Approximation and Projection (UMAP) as the primary methods for dimensionality reduction and visualization.
PCA is a widely used method for linear dimensionality reduction. PCA operates by finding a new orthogonal basis that captures directions of maximal variance in the data (shlens2014tutorialprincipalcomponentanalysis). By projecting high-dimensional vectors onto these principal axes, PCA allows one to rank-order components by their explanatory power. The method is computationally efficient, analytically grounded in linear algebra through eigenvalue decomposition or singular value decomposition, and does not require hyperparameter tuning. However, PCA is limited to capturing linear relationships and assumes that variance corresponds to meaningful structure, an assumption that may fail in non-Gaussian or highly nonlinear datasets (shlens2014tutorialprincipalcomponentanalysis). Despite these caveats, PCA remains a powerful baseline, is widely used in our experiments, and provides a stable reference frame for comparing geometric patterns.
On the other hand, Uniform Manifold Approximation and Projection (UMAP), proposed by McInnes et al. in 2018 (mcinnes2020umapuniformmanifoldapproximation), provides a nonlinear alternative designed to preserve both local neighborhoods and some aspects of global manifold structure. Rather than identifying variance-maximizing axes, UMAP builds a graph-based representation of data as a topological manifold and then optimizes a low-dimensional embedding that preserves this structure. In practice, UMAP is effective at revealing clustering patterns in latent states, producing interpretable visualizations. We chose UMAP over other nonlinear alternatives such as t-SNE (JMLR:v9:vandermaaten08a) because it better maintains global relationships while still revealing local clustering patterns. Its flexibility comes at the cost of tunable hyperparameters (e.g., number of neighbors, minimum distance) and a lack of linear structure, but it has become a widely adopted tool in machine learning for exploring latent structure.
@@ 2.4 Positional Embeddings/Encodings
In many early Transformer architectures, such as GPT-2 (used in our experiments), positional information is injected via learned positional embeddings. Each token position in the input sequence is associated with a learned vector that is added to its corresponding token embedding before entering the first Transformer block. Because these embeddings are learned during training, they often exhibit consistent geometric patterns that reflect the model’s internal representation of positional order. Within the context of our work, learned positional embeddings provide a straightforward baseline for the geometry of position-dependent features.
On the other hand, the LLaMa model we use employs a form of functional positional encoding called Rotary Positional Encodings (RoPE). Proposed by Su et al. in 2021 (su2023roformerenhancedtransformerrotary), RoPE applies a position-dependent rotation to query and key vectors within each attention head rather than adding fixed position vectors to the initial embeddings. This rotation encodes relative position information directly into the attention mechanism. Conceptually, RoPE embeds position in the geometry of relationships between tokens, rather than as an explicit position embedding vector.
This distinction serves as a key motivation behind Effects of Sequence Position experiments in section 4.5.
@@ 3.1 Pipeline
The pipeline supports two modes of operation. In the default “text” mode, text passages are sampled from a text dataset. For our experiments, we ran the model through the well-known Project Gutenberg (PG-19) dataset introduced by Rae et al. in 2019 (rae2019compressivetransformerslongrangesequence), consisting of a corpus of books written before 1919. Each passage is tokenized and truncated or padded to a fixed sequence length, ensuring uniformity across samples. In “singular” mode, by contrast, the pipeline samples individual tokens directly by iterating over the model’s vocabulary. This allows for fine-grained analysis of token-level representations without contextual interference. Both modes are parameterized by the number of samples and the desired sequence length (set at $1$ for the singular mode).
Figure 3 shows each of the capture points along the Transformer architecture. For each block, we record the outputs of each layer: both normalizations, the attention module, and the MLP. We denote the raw outputs of the attention and MLP components and their preceding norms as “pre-add” (points 1, 2, 4, and 5 in Figure 3), and the latent states after they have been added back into the residual stream as “post-add” (points 3 and 6 in Figure 3). In total, this yields six distinct captures per Transformer block.
@@ 3.2 Dimensionality Reduction and Visualization
Captured latent states are visualized in 22D following dimensionality reduction. Not all captured latent states must be used for dimensionality reduction and visualization. Instead, a subset may be selected (e.g., only latent states from particular layers). For dimensionality reduction, both PCA and UMAP are used, with a GPU-accelerated implementation for UMAP (via cuML (raschka2020machinelearningpythonmain)) utilized for visualizing large amounts of latent states. Before dimensionality reduction is performed, augmentations to the latent states can be applied. These include:
@@
  • Averaging the latent state vectors over any dimension(s) (sample dimension, sequence dimension, layer dimension). This may be done to remove the variance of a certain dimension to concentrate on the structure present in another.
PCA transformation is performed without mean centering to preserve the position of the origin.
We utilize a color-coding scheme to distinguish the layer and sequence dimensions. In our experiments, attention layers are blue while MLP layers are red. Darker shades indicate earlier layers while lighter shades indicate later ones. For the sequence dimension, we use a gradient to indicate earlier vs. later positions.
Fitted dimensionality reduction models can be reused; PCA fitted on one set of latent states can be applied to a different set of latent states, and visualizations can preserve x- and y-axis limits between runs. This is especially useful for ablation studies, where subtle differences between experimental conditions must be compared in a consistent coordinate frame. Additionally, dimensionality reduction can be performed that reduces the number of dimensions to a number greater than $2$. Then, any pair of the reduced dimensions can be visualized together in 22D (e.g., dimensions $2$ and $5$ or $3$ and $7$). Figure 11 shows the result of visualizing all possible pairs of dimensions after dimensionality reduction to $6$ dimensions (totaling ${6\choose 2}=15$ pairs), then stitching the resulting visualizations into a grid.
@@ 4.1 Experimental Setup
We use GPT-2 Large (774774M) and LLaMa-7B to generate the latent states. On all results using the PG-19 dataset, we use the following input shapes:
@@
We note that we use the maximum input length possible for both models. LLaMa is trained with a <BOS> (beginning of sentence) token prepended to all inputs. As a result, we prepend the <BOS> token to the beginning of all PG-19 inputs to LLaMa. After collecting the LLaMa latent states, we reduced the dimensionality of the latent states from $4096$ to $512$ using PCA. These dimensionality-reduced latent states were subsequently used in all visualizations. They were not used in any analyses of latent state norms; those used the original full dimensionality latent states for accurate results. All GPT-2 latent states were used in their full dimensionality.
Existing research demonstrates how different blocks in a Transformer have different roles. While earlier blocks focus on converting tokens into concepts as shown by Nanda et al. in 2023 (nanda2023factfinding) and the final blocks heavily denoise and refine the output, intermediate (middle) blocks, which are more robust to swapping and deletion, gradually create and refine intermediate features in a shared representation space as demonstrated by Sun et al. (sun2025transformerlayerspainters) and Lad et al. (lad2025remarkablerobustnessllmsstages) in 2024. For some of our experiments, we are interested in investigating the behavior of the intermediate block latent states without interference from earlier or later blocks. As a result, we define here the specific “intermediate” blocks/layers which we will later refer to in our results and discussions for both our tested models. Because we could not find a clear methodology to define concrete boundaries for the "intermediate" blocks, the definitions we make are somewhat arbitrary. Nevertheless, it is important for us to have consistent definitions across our analyses. For GPT-2, we define these as the layers in blocks 22-88 (from a range of 0-$11$), while for LLaMa, they are the layers in blocks 66-$27$ (from a range of 0-$31$).